[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Oct 05, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Oct 5 09:10:06 UTC 2016


HSU Shift Log for Wed Oct 05 2016 05:10:00 GMT-0400 (EDT)

Shift: A

Forecaster: D Brown

Technical Issues Summary:

ATCF issues: Percent symbol added into new Key Message #4 that cause an  
issue transmitting the TCD. Dave Zelinsky was at NHC and help fixed the  
problem. TCD went out a minute or two late.

N-AWIPS/AWIPS2 issues:

Other Techincal issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161005/9ce30d19/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list